OpenStudy (christos):

Improper integral can you show me how to solve this improper integral 1/ 1- x from 0 to 3

OpenStudy (christos):

I know it seems simple, I just want to understand something

OpenStudy (michele_laino):

I think that such integral is equivalent to these 2 independent limits: \[\Large \mathop {\lim }\limits_{{c_1} \to {1^ - }} \int_0^{{c_1}} {\frac{{dx}}{{x - 1}}} + \mathop {\lim }\limits_{{c_2} \to {1^ + }} \int_{{c_2}}^3 {\frac{{dx}}{{x - 1}}} \]

OpenStudy (christos):

thank you

OpenStudy (michele_laino):

:)